# Personhood credentials: Artificial intelligence and the value of

### v1

@@ 4.1   Equitable access
Benefits from personhood credentials must be weighed against potential impacts on the accessibility of digital services. A need for frequent authentication could contribute to friction and frustration for users, [^79] Although much of this paper focuses on the importance of preserving the Internet’s usability for people, many automated software scripts are also important for a well-functioning Internet. Such scripts could be inadvertently affected by the widespread introduction of PHCs; websites may want to consider approaches that offer benign bots a way to proactively distinguish themselves. One example of how some websites communicate their preferences for bot interaction is the use of robots.txt files, which have existed in some form since the 1990s as an opt-in approach for contending with the impacts of bots navigating the World Wide Web [149, 57]. particularly those less comfortable with technology generally (like older adults, who are common targets of AI-powered deception [246]). [^80] It is well recognized in commercial contexts that the friction of authentication is typically a negative experience for users [109, 114]. Depending on how authenticating with a PHC works, “Sign-in with PHC” options may ease some of this friction. If friction from PHCs is too high, some users may neglect to use them and thus lose out on the privacy benefits they afford.
@@
Relatedly, if service providers use personhood credentials to limit access to particular services, some groups of individuals without PHCs may be systematically excluded. It is thus important that PHC issuers design their systems with equitable access in mind, and that authorities guiding PHC ecosystem development are careful to ensure that all groups have some form of access to vital services. [^81] One analogous precedent may be found in the United States’s Blueprint for an AI Bill of Rights, which states that certain vital services should provide a way for users to opt out of AI interaction; see “Human Alternatives, Consideration, and Fallback” in [188]. Depending on the form of the PHC, some groups may have access before others: [^82] For example, if PHCs are primarily backed by government IDs, there will likely be geographic disparity in access to PHCs initially. Around 850 million people worldwide do not currently have an official identification document [272]. Refugee populations, in particular, may struggle to obtain or maintain a PHC. This issue must be investigated, perhaps under the purview of existing programs studying refugee documentation, e.g., within the United Nations [257]. It is important that this not contribute to the economic exclusion of slow-adopting groups in the meantime. [^83] PHCs will be less likely to contribute to economic exclusion if they are an opt-in mechanism for additional components of a service, rather than mandatory for the baseline service itself. One possibility is that PHCs may even help to unlock digital services for many who do not have access to forms of digital identity and are thus currently excluded from many basic digital services. [^84] It is estimated that 3.3 billion people worldwide do not currently have access to an official digital identity for online transactions—PHC systems could offer an easier-to-access alternative [272]. When considering PHCs’ impact on accessibility, it is important to note that as AI systems become increasingly indistinguishable from real people online, individuals and service providers may default to mistrusting digital strangers, which could also disproportionately affect some groups. These potential effects must likewise be assessed.
@@ 4.2   Free expression
It is important to consider how PHCs could reduce people’s willingness to speak freely and to dissent in digital spaces. Individuals may fear that their online speech will be linked to their offline identity, even though PHCs do not convey one’s identity. [^86] PHCs therefore respond to key criticisms of some platforms’ “real name” policies [30, 279]. Both perceived and actual anonymity are important for expressing views without fear of retribution, particularly for marginalized groups or for those living under authoritarian regimes with extensive surveillance [206, 153]. [^87] Anonymity has been referenced as an essential part of public discourse. The United States Supreme Court stated in McIntyre v. Ohio Elections Commission that anonymous speech “is not a pernicious, fraudulent practice, but an honorable tradition of advocacy and of dissent” [164]. See also [148]. In some cases, hostile governments may even be able to suppress authentic dissent by limiting access to PHCs within their borders. [^88] There are a number of ways that a hostile government might try to stop dissidents from getting PHCs. For instance, if a PHC system relies upon proprietary hardware, a hostile government might be able to restrict the hardware from its country, or otherwise control access to the hardware so that the government gains control over users’ PHCs [38]. Governments could also apply these prohibitions more narrowly to certain groups.
@@
PHCs also change the dynamics of credibility online, possibly in unpredictable ways that could affect individuals’ willingness to speak openly. Individuals who choose not to verify their personhood through PHCs might find their contributions discounted or even labeled as disinformation. [^89] This dynamic has some similarity to the concept of the liar’s dividend [225]. Though PHCs are implemented with the intent to mitigate online disinformation, they might have the side-effect of making non-PHC verified speech easier to write off as disinformation, even if it is in fact authentic. On the other hand, statements that might previously have been dismissed due to doubts about their origin could earn credibility when accompanied by PHC verification.
@@
While PHCs use unlinkable pseudonymity to preserve user privacy, they are not a remedy for pervasive surveillance practices like tracking and profiling used throughout the Internet today. [^90] Such practices are widely documented; for one comprehensive treatment, see [288]. Although PHCs prevent linking the credential across services, users should understand that their other online activities can still be tracked and potentially de-anonymized through existing methods [83]. It is important to recognize that if PHCs are challenging or inconvenient for users to use, companies may be inclined to direct them toward alternatives that are easier to use but provide less privacy. Consequently, some users might opt for these more user-friendly, yet less private, solutions. [^91] This is already seen with some examples of digital identity protocols and cryptocurrency wallets [146]. Emergent social dynamics may also threaten the anonymous nature of PHCs—for instance, if online abuse continues to be persistent even among individuals with PHCs, people may feel socially pressured to reveal auxiliary personal information to demonstrate that they are trustworthy.
@@
There are many nascent legal and ethical discussions centered on the rise of highly capable AI and freedom of expression. It will be crucial to understand how PHCs fit into the legal and ethical frameworks that emerge. For example, one important legal question is when individuals need to disclose their use of AI [143, 189, 258]. Service providers might wish to allow AI delegation on their platform only with proper transparency to other users, or might wish to disallow AI delegation on their service altogether. [^92] For example, a dating application may wish to disallow users delegating any conversations to an AI agent if they have a “verified person” account. Nonetheless, the service provider may have a difficult time determining that the user has illicitly delegated to an AI agent, if this delegation occurs after signing up using one’s PHC. If services wish to curtail delegation—particularly when users have incentives to illicitly delegate—services will need methods for detecting signals of AI-powered behavior (e.g., if a user is on the service at all hours of the day) and for enforcing rules against it thereafter. PHCs could be one component of disclosure rules for these new forms of expression, as well as a means of enforcing the rules that do emerge.
@@ 4.4   Robustness to attack and error
Beyond the attacks just described, a PHC system should also be robust to user error. For instance, PHC systems can offer methods to ensure that credentials can be recovered or revoked in case they are lost or stolen.103103103NIST offers recommendations on the management of cryptographic keys, which may be applicable to the design of recovery and revocation processes for PHCs [134]. The easier that it is for a holder to recover or revoke a compromised PHC, the less incentive an attacker will have to try to compromise a credential, as they should not expect to control the compromised PHC for long. On the other hand, if it is too easy to revoke one’s PHC and be reissued a new one, this may reduce the issuer’s ability to enforce its credential limit requirement of one per person per issuer. We discuss this challenge further in C Implementation choices for personhood credentials. A user should not be permanently locked out from vital digital services if they merely misplace their smartphone or other means of accessing their PHC; it is incumbent upon system designers to make PHCs as easy to use as possible for benign users.
@@ 5.1   Adapt existing digital systems
The emergence of highly capable AI technologies may render certain security assumptions inaccurate. Such assumptions may be common to a range of standards that governments and industry actors rely on. Key standards that may require reevaluation include the NIST’s Identity Assurance Levels (US),104104104NIST’s Digital Identity Guidelines are actively being revised [134, 181]. the European Union Agency for Cybersecurity’s (ENISA’s) practices for remote ID proofing,105105105ENISA’s 2024 report includes a full section on deepfake-powered attack vectors, though it also suggests certain indicators of deepfakes that we do not expect to persist—for instance, certain characteristics about their bitrate or frames per second [91]. and the international standard ISO/IEC 29115 [131].
Anti-spoofing recommendations are one component of identity standards that may now be outdated: Previously, asking a subject to replicate a certain pose on a live video call was reasonable evidence that they were a real person rather than a prerecorded video of one.106106106For instance, NIST’s guidance on Identity Verification suggests: “In order to confirm the video stream is live and not pre-recorded, the operator may direct the applicant to move their head in specific ways, raise or lower eyes, or ask the applicant questions requiring response during the live capture video” [180]. But malicious actors can increasingly use AI to mimic actions like these with lifelike avatars [147, 33, 276, 259, 123].
The relative strength of different forms of evidence may also deserve reevaluation based on AI’s improved abilities. For instance, authentication methods once considered effective in reducing fraud—such as voice-based authentication—are now increasingly vulnerable to AI-driven attacks. Advanced voice synthesis technologies can replicate an individual’s tone and speech patterns using minimal data, turning voice authentication into a potential avenue for account takeovers [196]. In some mediums (e.g., images and videos), the reliability of a piece of evidence might vary according to whether there is proof that the media was captured directly from a camera.107107107Photos taken with a C2PA-signing camera, like the smartphone application Truepic, can attest to being untampered [254]. More generally, reliability assessments of different forms of evidence may be improved by factoring in AI’s potential to subvert these forms of evidence.108108108For instance, NIST considers biometric verification to be its highest strength of evidence, even if done via remote identity proofing [180]. To remain confident in the strength of this evidence, standards bodies should consider explicitly stress-testing any biometric algorithms for robustness to deepfake attacks: As NIST notes, “The biometric False Match Rate (FMR)…does not account for spoofing attacks” [117].
@@
Currently, there are no standardized, official methodologies or tools to reliably measure the prevalence of AI-powered accounts across various digital platforms.109109109Without a standard approach, some have speculated that individual platforms have an incentive to permit some AI-powered accounts to drive up the total number of users and engagement, thus allowing the platforms to charge more money for advertising. For instance, ahead of Elon Musk’s acquisition of Twitter, he accused the site of engaging in these practices [69]. The company that Musk commissioned to conduct a study of Twitter’s bot activity—Cyabra—has claimed to have benchmarks across social media companies, but to our knowledge has not made these data or methodologies available [79]. Developing such tools could enable platforms and governments to perform robust comparative analysis and to determine the effectiveness of different mitigation strategies. Encouraging independent auditing firms to assess the presence of AI-generated content and accounts might promote transparent and unbiased analyses across communications platforms. Moreover, restoring or enhancing researchers’ access to platform data can offer deeper insights into the proliferation and trends of AI-driven accounts.110110110Some social media companies have phased out these forms of researcher access programs in recent years. In 2023, Twitter (now X) ended free access to APIs that had enabled research such as estimates of bot prevalence on the website [39]. More generally, some social media companies have taken action to restrict research programs studying dynamics on the sites [28].
@@ 5.2   Prioritize personhood credentials
  • How can users confirm that issuer and service provider privacy commitments are being upheld, without relevant cryptographic expertise?
Providing clear guidance on these issues will help stakeholders evaluate implementation trade-offs. Moreover, people can then engage in substantive dialogue that enhances PHCs through broad-based deliberation.
@@ Acknowledgments
We would like to thank the AI and Media Integrity Steering Committee of the Partnership on AI and the Berkman Klein Center for Internet & Society at Harvard University for hosting us for feedback sessions related to these topics. We would also like to thank a number of individuals for feedback and conversations: Scott Aaronson, Mada Aflak, Dan Alessandro, shirin anlen, Rahul Arora, James Aung, Boaz Barak, Katie Benjamin, Jake Brill, Miles Brundage, Jon Callas, Rosie Campbell, Alan Chan, Melissa Chase, Peter Cihon, Paul Crowley, Karen Easterbrook, Tyna Eloundou, Emiliano Falcon-Morano, Jason Fedor, Claudia Fischer, Tim Fist, Lydia Gorham, Margaret Hu, John Jordan, Gabe Kaptchuk, David Karger, Pedram Keyani, Grace Kwak Danciu, Kim Laine, Michael Lampe, Jaron Lanier, Teddy Lee, Brenda Leong, Leon Maksin, Timothy Marple, Reed McGinley-Stempel, Cullen O’Keefe, Christian Paquin, Joel Parish, Tobias Peyerl, Brendan Quinn, Ankur Rastogi, David Robinson, Ben Rossen, Girish Sastry, Bruce Schneier, Eric Scouten, Gaurav Sett, Yonadav Shavit, Adam Shostack, Allison Stanger, Jay Stanley, Esther Tetruashvily, Değer Turan, Raquel Vázquez Llorente, Chelsea Voss, Becky Waite, E. Glen Weyl, Rebecca Williams, and Jonathan Zittrain.
@@ B Relating personhood credentials to CAPTCHAs and synthetic content transparency tools
In this section, we further explore two common alternative approaches to the problem of scalable AI-powered deception online. PHCs generally either improve on or complement these approaches.

### v2

@@ 4.1   Equitable access
Benefits from personhood credentials must be weighed against potential impacts on the accessibility of digital services. A need for frequent authentication of one’s PHC could contribute to friction and frustration for users, [^79] Although much of this paper focuses on the importance of preserving the Internet’s usability for people, many automated software scripts are also important for a well-functioning Internet. Such scripts could be inadvertently affected by the widespread introduction of PHCs; websites may want to consider approaches that offer benign bots a way to proactively distinguish themselves. One example of how some websites communicate their preferences for bot interaction is the use of robots.txt files, which have existed in some form since the 1990s as an opt-in approach for contending with the impacts of bots navigating the World Wide Web [149, 57]. particularly those less comfortable with technology generally (like older adults, who are common targets of AI-powered deception [246]). [^80] It is well recognized in commercial contexts that the friction of authentication is typically a negative experience for users [109, 114]. Depending on how authenticating with a PHC works, “Sign-in with PHC” options may ease some of this friction. If friction from PHCs is too high, some users may neglect to use them and thus lose out on the range of benefits they afford.
@@
Relatedly, if service providers use personhood credentials to limit access to particular services, some groups of individuals without PHCs may be systematically excluded. It is thus important that PHC issuers design their systems with equitable access in mind, and that authorities guiding PHC ecosystem development are careful to ensure that all groups have some form of access to vital services. [^81] One analogous precedent may be found in the United States’s Blueprint for an AI Bill of Rights, which states that certain vital services should provide a way for users to opt out of AI interaction; see “Human Alternatives, Consideration, and Fallback” in [188]. Depending on the details of a PHC’s implementation, some groups may have access before others: [^82] For example, if PHCs are primarily backed by government IDs, there will likely be geographic disparity in access to PHCs initially. Around 850 million people worldwide do not currently have an official identification document [272]. Refugee populations, in particular, may struggle to obtain or maintain a PHC. This issue must be investigated, perhaps under the purview of existing programs studying refugee documentation, e.g., within the United Nations [257]. It is important that this not contribute to the economic exclusion of slow-adopting groups in the meantime. [^83] PHCs will be less likely to contribute to economic exclusion if they are an opt-in mechanism for additional components of a service, rather than mandatory for the baseline service itself. One positive possibility is that PHCs may help to unlock digital services for many who do not have access to forms of digital identity and are thus currently excluded from many basic digital services. [^84] It is estimated that 3.3 billion people worldwide do not currently have access to an official digital identity for online transactions—PHC systems could offer an easier-to-access alternative [272]. When assessing PHCs’ relative impact on accessibility, it is also important to consider that as AI systems become increasingly indistinguishable from real people online, individuals and service providers may default to mistrusting digital strangers, which could likewise disproportionately affect some groups.
@@ 4.2   Free expression
It is important to consider how PHCs could reduce people’s willingness to speak freely and to dissent in digital spaces. Individuals may fear that their online speech will be linked to their offline identity, even though PHCs do not convey one’s identity. [^86] PHCs therefore respond to key criticisms of some platforms’ “real name” policies [30, 279]. Both perceived and actual anonymity are important for expressing views without fear of retribution, particularly for marginalized groups or for those living under authoritarian regimes with extensive surveillance [206, 153]. [^87] Anonymity has been referenced as an essential part of public discourse. The United States Supreme Court stated in McIntyre v. Ohio Elections Commission that anonymous speech “is not a pernicious, fraudulent practice, but an honorable tradition of advocacy and of dissent” [164]. See also [148].
@@
PHCs also change the dynamics of credibility online, possibly in unpredictable ways that could affect individuals’ willingness to speak openly. Individuals who choose not to verify their personhood through PHCs might find their contributions discounted or even labeled as disinformation. [^88] This dynamic has some similarity to the concept of the liar’s dividend [225]. Though PHCs are implemented with the intent to mitigate online disinformation, they might have the side-effect of making non-PHC verified speech easier to write off as disinformation, even if it is in fact authentic. On the other hand, statements that might previously have been dismissed due to doubts about their origin could earn credibility when accompanied by PHC verification. In some cases, hostile governments might be able to suppress authentic and credible dissent by limiting access to PHCs within their borders. [^89] There are a number of ways that a hostile government might try to stop dissidents from getting PHCs. For instance, if a PHC system relies upon proprietary hardware, a hostile government might be able to restrict the hardware from its country, or otherwise control access to the hardware so that the government gains control over users’ PHCs [38]. Governments could also apply these prohibitions more narrowly to certain groups.
@@
While PHCs preserve user privacy via unlinkable pseudonymity, they are not a remedy for pervasive surveillance practices like tracking and profiling used throughout the Internet today. [^90] Such practices are widely documented; for one comprehensive treatment, see [288]. Although PHCs prevent linking the credential across services, users should understand that their other online activities can still be tracked and potentially de-anonymized through existing methods [83]. It is important to recognize that if PHCs are challenging or inconvenient for users to use, companies may be inclined to direct them toward alternatives that are easier to use but provide less privacy. Consequently, some users might opt for these more user-friendly, yet less private, solutions. [^91] This is already seen with some examples of digital identity protocols and cryptocurrency wallets [146]. Emergent social dynamics may also threaten the anonymous nature of PHCs—for instance, if online abuse continues to be persistent even among individuals with PHCs, people may feel socially pressured to reveal auxiliary personal information to demonstrate that they are trustworthy.
@@
There are many nascent legal and ethical discussions centered on the rise of highly capable AI and freedom of expression. It will be crucial to understand how PHCs fit into the legal and ethical frameworks that emerge. For example, one important legal question is when individuals need to disclose their use of AI [143, 189, 258]. Service providers might wish to allow AI delegation on their platform only with proper transparency to other users, or they might wish to disallow AI delegation on their service altogether. [^92] For example, a dating application may wish to disallow users delegating any conversations to an AI agent if they have a “verified person” account. Nonetheless, the service provider may have a difficult time determining that the user has illicitly delegated to an AI agent, if this delegation occurs after signing up using one’s PHC. If services wish to curtail delegation—particularly when users have incentives to illicitly delegate—services will need methods for detecting signals of AI-powered behavior (e.g., if a user is on the service at all hours of the day) and for enforcing rules against it thereafter. PHCs could provide one method for disclosing these new forms of expression, as well as a means of enforcing the related rules that do emerge.
@@ 4.4   Robustness to attack and error
Beyond the attacks just described, a PHC system should also be robust to user error. For instance, PHC systems can offer methods to ensure that credentials can be recovered or revoked in case they are lost or stolen.103103103NIST offers recommendations on the management of cryptographic keys, which may be applicable to the design of recovery and revocation processes for PHCs [134]. The easier that it is for a holder to recover or revoke a compromised PHC, the less incentive an attacker will have to try to compromise a credential, as they should not expect to control the compromised PHC for long. On the other hand, if it is too easy to revoke one’s PHC and be reissued a new one, this may reduce the issuer’s ability to enforce its credential limit requirement of one per person per issuer. We discuss this challenge further in C Implementation choices for personhood credentials. A user should not be permanently locked out from vital digital services if they merely misplace their smartphone or other means of accessing their PHC; it is incumbent upon system designers to make PHCs error-tolerant for benign users.
@@ 5.1   Adapt existing digital systems
The emergence of highly capable AI technologies may render certain security assumptions inaccurate. Such assumptions may be common to a range of standards that governments and industry actors rely on. Key standards that may require reevaluation include NIST’s Identity Assurance Levels (US),104104104NIST’s Digital Identity Guidelines are actively being revised [134, 181]. the European Union Agency for Cybersecurity’s (ENISA’s) practices for remote ID proofing,105105105ENISA’s 2024 report includes a full section on deepfake-powered attack vectors, though it also suggests certain indicators of deepfakes that we do not expect to persist—for instance, certain characteristics about their bitrate or frames per second [91]. and the international standard ISO/IEC 29115 regarding entity authentication [131].
Anti-spoofing recommendations are one component of standards that may now be outdated: Previously, asking a subject to replicate a certain pose on a live video call was reasonable evidence that they were a real person rather than a prerecorded video of one.106106106For instance, NIST’s guidance on Identity Verification suggests: “In order to confirm the video stream is live and not pre-recorded, the operator may direct the applicant to move their head in specific ways, raise or lower eyes, or ask the applicant questions requiring response during the live capture video” [180]. But malicious actors can increasingly use AI to mimic actions like these with lifelike avatars [147, 33, 276, 259, 123].
The relative strength of different forms of evidence may also warrant reevaluation based on AI’s improved abilities. For instance, authentication methods once considered effective in reducing fraud—such as voice-based authentication—are now increasingly vulnerable to AI-driven attacks. Advanced voice synthesis technologies can replicate an individual’s tone and speech patterns using minimal data, turning voice authentication into a potential avenue for account takeovers [196]. In some domains (e.g., images and videos), the reliability of a piece of evidence might vary according to whether there is proof that the media was captured directly from a camera.107107107Photos taken with a C2PA-signing camera, like the smartphone application Truepic, can attest to being untampered [254]. More generally, reliability assessments of different forms of evidence may be improved by directly factoring in how AI might be used to subvert the evidence.108108108For instance, NIST considers biometric verification to be its highest strength of evidence, even if done via remote identity proofing [180]. To remain confident in the strength of this evidence, standards bodies should consider explicitly stress-testing any biometric algorithms for robustness to deepfake attacks: As NIST notes, “The biometric False Match Rate (FMR)…does not account for spoofing attacks” [117].
@@
Currently, there are no standardized, official methodologies or tools to reliably measure the prevalence of AI-powered accounts across various digital platforms.109109109Without a standard approach, some have speculated that individual platforms have an incentive to permit some AI-powered accounts to drive up the total number of users and engagement, thus allowing the platforms to charge more money for advertising. For instance, ahead of Elon Musk’s acquisition of Twitter, he accused the site of engaging in these practices [69]. The company that Musk commissioned to conduct a study of Twitter’s bot activity—Cyabra—has claimed to have benchmarks across social media companies, but to our knowledge has not made these data or methodologies available [79]. Developing such tools could enable platforms and governments to perform robust comparative analysis and to determine the effectiveness of different mitigation strategies. Encouraging independent auditing firms to assess the presence of AI-generated content and accounts might promote transparent and unbiased analyses across communications platforms. Moreover, restoring or enhancing researchers’ access to platform data can offer insights into the proliferation and trends of AI-driven accounts.110110110Some social media companies have phased out these forms of researcher access programs in recent years. In 2023, Twitter (now X) ended free access to APIs that had enabled research such as estimates of bot prevalence on the website [39]. More generally, some social media companies have taken action to restrict research programs studying dynamics on the sites [28].
@@ 5.2   Prioritize personhood credentials
  • How can users confirm that privacy commitments from an issuer and service provider are being upheld, without relevant cryptographic expertise?
Research that explores these issues will help stakeholders evaluate implementation trade-offs. Moreover, people can then engage in substantive dialogue that enhances PHCs through broad-based deliberation.
@@ Acknowledgments
We would like to thank the AI and Media Integrity Steering Committee of the Partnership on AI and the Berkman Klein Center for Internet & Society at Harvard University for hosting us for feedback sessions related to these topics. We would also like to thank a number of individuals for feedback and conversations: Scott Aaronson, Mada Aflak, Dan Alessandro, shirin anlen, Rahul Arora, James Aung, Boaz Barak, Katie Benjamin, Jake Brill, Miles Brundage, Jon Callas, Rosie Campbell, Alan Chan, Melissa Chase, Peter Cihon, Paul Crowley, Karen Easterbrook, Tyna Eloundou, Emiliano Falcon-Morano, Jason Fedor, Claudia Fischer, Tim Fist, Lydia Gorham, Mark Gray, Margaret Hu, John Jordan, Gabe Kaptchuk, David Karger, Pedram Keyani, David Kim, Grace Kwak Danciu, Kim Laine, Michael Lampe, Jaron Lanier, Teddy Lee, Traci Lee, Brenda Leong, Leon Maksin, Timothy Marple, Stephen McAleer, Reed McGinley-Stempel, Pamela Mishkin, Ben Newhouse, Cullen O’Keefe, Christian Paquin, Joel Parish, Tobias Peyerl, Brendan Quinn, Ankur Rastogi, Francis Real, Elizabeth M. Renieris, David Robinson, Ben Rossen, Girish Sastry, Bruce Schneier, Eric Scouten, Gaurav Sett, Yonadav Shavit, Dane Sherburn, Adam Shostack, Allison Stanger, Jay Stanley, Esther Tetruashvily, Değer Turan, Raquel Vázquez Llorente, Chelsea Voss, Becky Waite, E. Glen Weyl, Rebecca Williams, Wenda Zhou, and Jonathan Zittrain.
@@ B Relating personhood credentials to CAPTCHAs and synthetic content transparency tools
In this section, we further explore two common alternative approaches to the problem of scalable AI-powered deception online. PHCs can complement these approaches.